If Beth can walk two and a half miles in 28 minutes how many minutes will it take her to walk 10 miles
Answer by rapture(86)

Set up a proportion.

Let m = number of minutes we need to find.

2_1/2 miles = 5/2 miles

Here is the set up:


(5/2)/(10) = (28/m)


(10)(28) = (5/2)(m)


280 = (5m/2)

Solve for m to find your answer.